OpenStudy (donnie1999):

Frank has devised a formula for his catering business that calculates the number of meatballs he needs to prepare. The formula is m= 4a+2c, where c= number of children; m= number of meatballs; a= number of adults How many meat balls are required for a party of 20 adult and 8 children?

OpenStudy (donnie1999):

@crissyyyxoxo

OpenStudy (anonymous):

m= 4(25) + 2(6) = 100 + 12 = 112

OpenStudy (anonymous):

110

OpenStudy (donnie1999):

it was wrong:(

OpenStudy (anonymous):

did you try 112?